# Shoal++: High Throughput DAG BFT Can Be Fast!

### v1

@@ 8.1. Failure-free case
Figure 5 illustrates throughput and latency under failure-free conditions. Shoal++ is the only system to sustain sub-second e2e latency up to a throughput of 100k transactions per second (tps) while simultaneously scaling throughput to over 150k tps. At low load, Shoal++ achieves 775ms median latency.
Jolteon is bottlenecked by the leader’s network bandwidth and cannot scale throughput beyond 2100 tps. At a low load, Jolteon has 900ms median latency.
@@
Bullshark and Shoal support much higher throughput but suffer from high latency. They incur, even at  1400 tps, respectively 1.9s and 1.45s median e2e latency. With 50k tps the latency increases to 2.4 and 1.7, respectively. This follows directly from their high average number of required message delays (§3.2). The throughput of Bullshark and Shoal is bottlenecked by network utilization since we disabled Narwhal streaming for better latency.
@@
Shoal++ significantly reduces latency (at least by 50% over Shoal, and 60% over Bullshark), while simultaneously offering improved throughput. The parallel DAG construction allows Shoal++ to better utilize network and storage resources: rather than broadcasting and processing a large batch once per round, it broadcasts and processes three smaller batches at the same time.
Mysticeti matches the throughput of Shoal++, which suggests that they both reach a networking bandwidth bottleneck. However, Mysticeti fails to match the latency of Shoal++ (and Jolteon at low load). This is presumably because at higher loads – even in the absence of faults – replicas may be temporarily out of sync and need to fetch missing data on the critical path of consensus, resulting in a slowdown.
Figure 6. Shoal++ Breakdown. No failures with 100 geo-distributed replicas.
8.2. Latency improvement breakdown
@@
Figure 6 isolates the respective latency improvement of each Shoal++ augmentation compared to the baseline, Shoal. Shoal++ Faster Anchors shows Shoal augmented only with the Fast Direct Commit Rule. Shoal++ More Faster Anchors adds the multiple anchor augmentation to the former, and omits the parallel DAGs optimization. The Fast Commit rule in theory improves anchor commit latency by 2md. In real, asymmetric network settings, however, it is not always 2md faster than the Direct Commit rule, and thus the practical latency improvement is smaller.
@@
The multiple anchor augmentation contributes to improved latency the most. First, it saves on average 3md by eliminating the anchoring latency for most nodes. Second, it makes sure that the anchoring latency for non-anchor nodes is one DAG round. This is because, with many anchors per round, each non-anchor node is referenced by at least one anchor with an overwhelming probability. The parallel DAG framework improves latency further by reducing queuing latency, but more importantly, it improves throughput scalability by approaching a the data ”streaming” effect; crucially, and unlike Narwhal-based worker designs, it does so without incurring additional latency.
Figure 7. Latency/Throughput graph with 33 out of 100 crashed replicas.
8.3. Failure case
To understand how robust Shoal++ is, we compared it to the baseline systems under less favorable conditions with 33 crash failures out of a total number of 100 replicas. The results are shown in Figure 7. Unfortunately, we could not run Mysticeti in this scenario because distributing failures across regions is not supported in their codebase. [^8] Mysticeti allows only for specific regions to fail collectively. Jolteon demonstrates roughly the same latency under low load as their leader reputation mechanism quickly overcomes the failures and elects only alive replicas. Bullshark, in contrast, suffers from higher latency and lower TP since it does not have a leader reputation mechanism to elect the strongest anchors. Shoal and Bolt keep roughly the same throughput as in the failure-free case but suffer from higher latency due to a higher quorum time - since failures are equally distributed, a quorum of replicas has to span more regions.
9. Related work
@@
Spurred by the seminal PBFT (Castro and Liskov, 2002) protocol, a long line of systems (Kotla et al., 2007; Gueta et al., 2019; Martin and Alvisi, 2006; Suri-Payer et al., 2021; Gelashvili et al., 2022) optimize for consensus latency. PBFT-based protocols designate a single leader to act as sequencer, and replace it only sporadically (e.g upon failure, or eagerly for fairness). This approach enjoys excellent latency. PBFT requires, when led by a correct replica, only 3 message delays (md); the optimum achievable latency in presence of faults, and optimal resilience of $n=3f+1$. Several followup works suggest optimistic Fast Paths (Kotla et al., 2007; Gueta et al., 2019; Martin and Alvisi, 2006; Suri-Payer et al., 2021) that improve latency to only 2md in absence of faults, or by weakening resilience to $n=5f+1$. Unfortunately, single leader-based designs are fundamentally bottlenecked in throughput by the bandwidth and processing capacity of a single replica.
@@
The key idea to scale beyond this bottleneck was realizing that networking resources must be fully utilized. Some works propose to decouple data dissemination from the consensus logic (Danezis et al., 2022; Yang et al., 2022; Arun and Ravindran, 2022), while others (Stathakopoulou et al., 2019; Gupta et al., 2021; Stathakopoulou et al., 2022) propose multi-log frameworks: these systems partition the request space and operate multiple parallel black-box instances of consensus protocols (e.g. PBFT) – each led by a different replica –, and carefully intertwine their outputs to construct a single, totally ordered log. This approach can achieve high throughput but requires complex coordination to deal with failures and re-configuration.
@@
[…]
Narwhal and Tusk (Danezis et al., 2022) catapulted DAG-BFT to popularity by demonstrating that (i) (asynchronous) consensus can be embedded onto DAG-structures without the need for explicit view changes (Tusk), and (ii) throughput can be scaled nearly linearly through the addition of a horizontally scalable worker layer (Narwahl Mempool). Bullshark (Spiegelman et al., 2022a) improved upon Tusk by adopting partial synchrony and introducing a common case fast path during synchronous intervals; in a future iteration (Spiegelman et al., 2022b) Bullshark removed the asynchronous fallback, and established itself as the de facto partially synchronous DAG-BFT protocol.
@@
A recent group of protocols propose the design of uncertified DAGs. Such constructions can, in theory, reduce latency, but are increasingly brittle as they are prone to data fetching on the critical path (§3.2). Cordial Miners (Keidar et al., 2022) replaces the Reliable Broadcast (RB) in each round with best-effort broadcast (BEB), allowing anchors to (in the best case) commit within 3md; anchors are placed every other round, for an average latency of 4.5md. Similarly, BBCA-chain (Malkhi et al., 2023) replaces non-anchor nodes with BEB, and strengthens anchor nodes to single-shot pbft instances. Unfortunately, neither proposal has an implementation that empirically evaluates practical performance.
Concurrent Latency reduction efforts.
@@
Sailfish (Shrestha et al., 2024) introduces a new anchor commit rule that resembles the Fast Direct Commit Rule proposed in  Shoal++, but unlike Shoal++ forgoes leveraging the Bullshark Direct Commit rule as potentially faster alternative. Unfortunately, Sailfish lacks an implementation that empirically substantiates their suggested latency improvement.
Mysticeti improves upon Bullshark’s best case latency by transitioning to an uncertified DAG and implementing the Cordial Miners (Keidar et al., 2022) consensus protocol. Mysticeti, like Shoal++, adopts Shoal’s high-level proposal to optimistically employ multiple anchors per round. Our evaluation demonstrates that Mysticeti almost matches Shoal++’s latency in the common case. However, as discussed in §3.2, and independently corroborated by Motorway (Giridharan et al., 2024), uncertified DAG constructions are prone to data fetching on the critical path, which can, in practice, negate the suggested latency benefits. A single Byzantine (or slow) replica, for instance, can impose at least 2 additional message delays per round in Mysticeti by not fully (or timely) disseminating its node proposals. The certified DAG employed by Shoal++ significantly enhances its resilience to Byzantine behavior.
Motorway (Giridharan et al., 2024) introduces a promising DAG-free consensus approach in which replicas efficiently and reliably disseminate data independently in parallel ”data lanes”, and consensus commits only snapshot cuts of the lane states.
@@
To the best of our knowledge, leader reputation for BFT systems was first implemented in DiemBFT (Team, 2021), and formalized by Carousel (Cohen et al., 2022). Shoal (Spiegelman et al., 2023) and later Hammerhead (Tsimos et al., 2023) adopt leader reputation to DAG-BFT in an effort to exclude underperforming leaders from anchor candidacy. Shoal++ extends the approach to provide, in each round, a set of anchor candidates.
10. Conclusion
This work introduces Shoal++, a DAG-based BFT consensus protocol that offers high throughput, while significantly lowering e2e latency compared to current state of the art DAG-BFT protocols. Shoal++ takes as its starting point existing state of the art DAG-BFT protocols, and augments them in three key ways: (i) it reduces, in the common case, the commit latency of leaders (anchors), (ii) dynamically reinterprets anchor schedules to increase anchor frequency, and (iii) operates multiple parallel DAG instances to increase proposal frequency. This, in sum, allows Shoal++ to reduce the fault-free average e2e latency from 10.5 message exchanges (Shoal (Spiegelman et al., 2023)) to only 4.5 message exchanges, within reach of traditional latency optimal (but throughput-inefficient) BFT protocols such as PBFT.

### v2

@@ 8.1 Failure-free case
Figure 5 illustrates throughput and latency under failure-free conditions. Shoal++ is the only system to sustain sub-second latency for throughput of 100k transactions per second (tps).
@@
Jolteon offers low latency (900ms) at low load, but is unable to scale throughput beyond 2100 tps as it is bottlenecked by the leader’s network bandwidth. Bullshark and Shoal support significantly higher throughput (up to 75k tps) but incur high latency: 1.9s and 1.45s, respectively, at low load, and 2.4s and 1.7s, respectively at 50k tps. This follows directly from their high average number of required message delays (§3.2). Throughput scalability in both systems is bottlenecked by network utilization.
@@
Shoal++ significantly reduces latency (775ms at low load), while simultaneously improving throughput scalability (up to ca. 140k tps). The parallel DAG construction allows Shoal++ to better utilize network and storage resources: rather than broadcasting and processing a large batch once per round, it broadcasts and processes three smaller batches at the same time. Applying the same technique to Bullshark and Shoal ("More DAGs") allows both systems to match the throughput of Shoal++ (while simultaneously improving queuing latency), without incurring the latency penalty of a Narwahl-style worker layer.
Mysticeti matches the throughput of Shoal++, but fails to match its latency, even at moderate loads. At high loads (upwards of 100k tps) Mysticeti notably incurs a signficant latency penalty as replicas may be temporarily out of sync, and must fetch missing data on the critical path of consensus.
Figure 6: Shoal++ Breakdown. No failures with 100 geo-distributed replicas.
Figure 7: Latency/Throughput graph with 33 out of 100 crashed replicas.
8.2 Latency improvement breakdown
@@
Figure 6 isolates the respective latency improvement of each Shoal++ augmentation compared to the baseline, Shoal. Shoal++ Faster Anchors shows Shoal augmented only with the Fast Direct Commit Rule while Shoal++ More Faster Anchors additionally applies the multiple anchor augmentation. Finally, Shoal++ includes also the parallel DAG optimization.
@@
The Fast Commit rule, in theory, improves anchor commit latency by up to 2md. In real, asymmetric network settings, however, it is not always 2md faster than the Direct Commit rule, and thus the practical latency improvement is smaller. The multiple anchor augmentation, in turn, improves latency significantly: (i) it saves, on average, 3md by eliminating the anchoring latency for most nodes (those that become anchors), and (ii) ensures that the anchoring latency for non-anchor nodes is typically only one DAG round because (with overwhelming probability) nodes are referenced by at least one of the (many) anchors in the next round. The parallel DAG framework improves latency further, by reducing queuing latency, but more importantly improves throughput scalability by approaching a data "streaming" effect; crucially, and unlike Narwhal-based worker designs, it does so without incurring additional latency.
8.3 Failure case
To quantify the robustness of Shoal++, relative to the baseline systems, we evaluate performance under two types of disruptions.
First, we measure the performance of all systems while simulating crash failures of 33 (out of 100) replicas (Fig. 7). The performance of Jolteon remains largely unaffected, as its leader reputation mechanism quickly detects failures and subsequently elects only alive replicas. Shoal and Shoal++, likewise, swiftly adjust and elect only live anchors; latency, however, increases (up to ca. 2x at high load) as quorums must span more regions, increasing tail latency. Bullshark and Mysticeti, in contrast, suffer drastically increased latency since they do not employ a leader reputation mechanism to elect suitable anchor candidates, and thus must "wait" to commit until a non-faulty replica is elected as anchor.
Second, we measure the impact of sporadic message drops for Shoal++ and Mysticeti (Fig. 8). We inject network layer message drops for 1% of egress traffic in 5 nodes (out of 100 total), beginning at 60 seconds (red line). Under normal network conditions and a low-moderate load (18k tps) both Shoal++ and Mysticeti offer 700ms median consensus latency. Upon failure injection the latency of Mysticeti rises sharply (by a factor of 10x) as replicas scramble to perform critical-path synchronization on missing data (resulting, at times, in timeouts). As latency spikes, throughput initially drops as well, but recovers once the backlog of disseminated DAG proposals commit. Shoal++, in contrast, remains largely unaffected by message drops (latency rises to at most 1.3x). Because all nodes are certified, the DAG construction process in Shoal++ proceeds smoothly; any required synchronization is asynchronous and off the critical path.
Figure 8: Impact of message drops in Certified vs Uncertified DAGs.
9 Related work
@@
Spurred by the seminal PBFT [14] protocol, a long line of systems [27, 23, 29, 38, 21] optimize for consensus latency. PBFT-based protocols designate a single leader to act as sequencer, and replace it only sporadically (e.g upon failure, or eagerly for fairness). This approach enjoys excellent latency. PBFT requires, when led by a correct replica, only 3 message delays (md); the optimum achievable latency in presence of faults, and optimal resilience of $n=3f+1$. Several followup works suggest optimistic Fast Paths [27, 23, 29, 38] that improve latency to only 2md in absence of faults, or by weakening resilience to $n=5f+1$. Unfortunately, single leader-based designs are fundamentally bottlenecked in throughput by the bandwidth and processing capacity of a single replica.
@@
The key idea to scale beyond this bottleneck was realizing that networking resources must be fully utilized. Some works propose to decouple data dissemination from the consensus logic [18, 45, 10], while others [36, 24, 37] propose multi-log frameworks: these systems partition the request space and operate multiple parallel black-box instances of consensus protocols (e.g. PBFT) – each led by a different replica –, and carefully intertwine their outputs to construct a single, totally ordered log. This approach can achieve high throughput but requires complex coordination to deal with failures and re-configuration.
@@
HashGraph [12] was the first to propose a DAG-based consensus structure, and Danezis et al. [31] showed that any deterministic BFT protocol can, in theory, be embedded on a DAG. Aleph [20] and DAG-Rider [25] were the first to introduce additional rigidity to the DAG construction, leveraging a round-based structure to design asynchronous BFT protocols. Blockmania [17] was the first attempt to implement a DAG-BFT system, but fell short of adoption as it inherited a complex view change procedure, notorious to traditional BFT protocols such as PBFT.
Narwhal and Tusk [18] catapulted DAG-BFT to popularity by demonstrating that (i) (asynchronous) consensus can be embedded onto DAG-structures without the need for explicit view changes (Tusk), and (ii) throughput can be scaled nearly linearly through the addition of a horizontally scalable worker layer (Narwahl Mempool). Bullshark [34] improved upon Tusk by adopting partial synchrony and introducing a common case fast path during synchronous intervals; in a future iteration [35] Bullshark removed the asynchronous fallback, and established itself as the de facto partially synchronous DAG-BFT protocol.
@@
A recent group of protocols propose the design of uncertified DAGs. Such constructions can, in theory, reduce latency, but are increasingly brittle as they are prone to data fetching on the critical path (§3.2). Cordial Miners [26] replaces the Reliable Broadcast (RB) in each round with best-effort broadcast (BEB), allowing anchors to (in the best case) commit within 3md; anchors are placed every other round, for an average latency of 4.5md. Similarly, BBCA-chain [28] replaces non-anchor nodes with BEB, and strengthens anchor nodes to single-shot pbft instances. Unfortunately, neither proposal has an implementation that empirically evaluates practical performance.
Concurrent Latency reduction efforts
@@
Sailfish [32] introduces a new anchor commit rule that resembles the Fast Direct Commit Rule proposed in  Shoal++, but unlike Shoal++ forgoes leveraging the Bullshark Direct Commit rule as potentially faster alternative. Unfortunately, Sailfish lacks an implementation that empirically substantiates their suggested latency improvement.
Mysticeti improves upon Bullshark’s best case latency by transitioning to an uncertified DAG and implementing the Cordial Miners [26] consensus protocol. Mysticeti, like Shoal++, adopts Shoals high-level proposal to optimistically employ multiple anchors per round. Our evaluation demonstrates that Mysticeti almost matches Shoal++’s latency in the common case. However, as shown in §8, and independently corroborated by Autobahn [22], uncertified DAG constructions are prone to data fetching on the critical path, which negates the suggested latency benefits. A single Byzantine (or slow) replica, for instance, can impose at least 2 additional message delays per round in Mysticeti by not fully (or timely) disseminating its node proposals. In §8 we observed 10x latency degradation when 5 out of 100 nodes experienced 1% message drop.
Autobahn [22] introduces a promising DAG-free consensus approach in which replicas efficiently and reliably disseminate data independently in parallel "data lanes", and consensus commits only snapshot cuts of the lane states.
@@
To the best of our knowledge, leader reputation for BFT systems was first implemented in DiemBFT [42], and formalized by Carousel [16]. Shoal [33] and later Hammerhead [44] adopt leader reputation to DAG-BFT in an effort to exclude underperforming leaders from anchor candidacy. Shoal++ extends the approach to provide, in each round, a set of anchor candidates.
10 Conclusion
This work introduces Shoal++, a DAG-based BFT consensus protocol that offers high throughput, while significantly lowering e2e latency compared to current state of the art DAG-BFT protocols. Shoal++ takes as its starting point existing state of the art DAG-BFT protocols, and augments them in three key ways: (i) it reduces, in the common case, the commit latency of leaders (anchors), (ii) dynamically reinterprets anchor schedules to increase anchor frequency, and (iii) operates multiple parallel DAG instances to increase proposal frequency. This, in sum, allows Shoal++ to reduce the fault-free average e2e latency from 10.5 message exchanges (Shoal [33]) to only 4.5 message exchanges, within reach of traditional latency optimal (but throughput-inefficient) BFT protocols such as PBFT.